OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2x+4 (_,_) Please Help!!

OpenStudy (anonymous):

Just substitute the first equation into the second and solve for x: \[ \frac {3}{2} x = - \frac {1}{2} x + 4 \implies 2x = 4 \implies x = 2 \]Now, since y = 3/2 * x, we have y = 3 and we are done.

OpenStudy (gorv):

is it \[y=\frac{ 3 }{ 2 }x\]

OpenStudy (anonymous):

You can put this in slope intercept form, do you know how to do that?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

Yeah it is Y=3/2x

OpenStudy (anonymous):

so is the answer 4,2 because that is what I was thinking but I don't know

OpenStudy (anonymous):

never mind the answer is (2,3)

OpenStudy (gorv):

u have y=3/2x replace y in other equation